Find the effective rate of the compound interest rate or investment. (Round your answer to two decimal places.)

A $90,000 zero-coupon bond maturing in 9 years and selling now for $43,035.
Answer #1
Face value 90000
Current market price 43035
Time in years 9
Face value= Current Price * (1+Interest rate)^Time
90000= 43035*(1+Interest rate)^9
90000/43035= (1+Interest rate)^9
2.091321018 (1+Interest rate)^9
2.0913^(1/9)= (1+Interest rate)
1.085429986 (1+Interest rate)
Interest rate 0.08543
Interest rate 8.54%